OpenStudy (anonymous):

Check all polynomials that are perfect square trinomials: 49x2 − 8x + 16 4a2 − 20a + 25 25b2 − 5b + 10 16x2 − 8x + 1

OpenStudy (anonymous):

B . D . ;)

OpenStudy (anonymous):

can u explain

OpenStudy (anonymous):

4a² − 20a + 25: discriminant = (-20)² - 4(4)(25) = 0 16x² − 8x + 1: discriminant = (-8)² - 4(16)(1) = 0 The other two aren't these .

OpenStudy (anonymous):

Thanks ! ;)

OpenStudy (anonymous):

can u help with two more ? i have to put the answers exactly right or i get it wrong

OpenStudy (anonymous):

Sure as long as there not as hard as the other one was . xD

OpenStudy (anonymous):

Given the following perfect square trinomial, fill in the missing term. x2 − 16x + ____

OpenStudy (anonymous):

even if i have the right answer, if i put in one wrong anything anywhere it will be marked as wrong :/

OpenStudy (anonymous):

(x−8)2=x2−16x+64 . How this ?

OpenStudy (anonymous):

Given the following perfect square trinomial, fill in the missing term. (Do not type the variable in the blank.) 4x2 + ___x + 49 idk ill have to see when i submit it

OpenStudy (anonymous):

wait is that teh answer itself or equation = answer

OpenStudy (anonymous):

Ummm , I don't know what you mean but , the answer for the first question you asked was 64 .

OpenStudy (anonymous):

:/ crap. well it came up wrong :( so instead of an A i got a D. i hate those questions

OpenStudy (anonymous):

Noo ! i didn't answer the last one yet . It was 28 .

OpenStudy (anonymous):

Sorry , if I gave you a low grade .

OpenStudy (anonymous):

yea i figured out the last one lol and its not your fault, its the stupid database

OpenStudy (anonymous):

Well , I'll just give you a medal since I feel guilty .

OpenStudy (anonymous):

aw well thanks for helping.

OpenStudy (anonymous):

Your Welcome .